In March 2012, Daniela motor financing, offered some securities for sale to the public. Under the terms of the deal, DMF promise to repay the owner of one of the securities $500 in March 2037, but investors would receive nothing until then. Investors PDF to earn $50 for each of the securities; so they gave up $250 in March 2012, for the promise of a $500 payment 25 years later.
c. In 2022, instead of cashing in the van for it then current value, you decide to hold the bond until it matures in 2037. What annual rate of return will you earn over the last 15 years?
Answer #1

a) Initial return (500/250)^(1/25)-1 = 2.81%

b) Worth after 10 years at 0.8% (250*(1+0.8%)^10) = 270.74

c) Return in last 15 years (500/270.74)^(1/15)-1 = 4.17%